Given the function f(x)=-x^3/3+x^2-6x-2, discuss its relative maximum and minimum

points, the intervals where it is increasing and decreasing, the intervals of concavity, and the points of inflection. Construct a sketch of the graph of the function.



1
Expert's answer
2021-09-27T23:27:30-0400

Given the function "f(x)=-\frac{x^3}3+x^2-6x-2", let us discuss its relative maximum and minimum

points, the intervals where it is increasing and decreasing, the intervals of concavity, and the points of inflection.

Since "f'(x)=-x^2+2x-6=-(x^2-2x+1)-5=-(x-1)^2-5<0," we conclude that the function is decreasing on the set "\R" of real numbers. It has no relative maximum and minimum points.

Taking into account that "f''(x)=-2x+2," and "f''(x)=0" implies "x=1," we conclude that "x=1" is the point of inflection. Since "f''(x)>0" for "x<1," we get that on the interval "(-\infty,1)" the function is concave up. Since "f''(x)<0" for "x>1," we conclude that on the interval "(1,+\infty)" the function is concave down.


Let us construct a sketch of the graph of the function: